What is the 100th term of series 1+3+6+10+...........

Grade:11

1 Answers

Vinod Prajapati
6 years ago
By observation the given series is in the following way:For 1st term 1(1+1)/2=1For 2nd term2(2+1)/2=3For 3rd term 3(3+1)/2=6And so on,Therefore for nth term we can write asn(n+1)/2So for 100 th term100(100+1)/2=5050